3 charges are positioned along a horizontal line.

A +5 C charge is placed at the point X = 0 m

A +7 C charge is placed at the point X = 0.2 m

A -1 C charge is placed at the point X = 1.4 m.

What is the magnitude of the force on the +5 C charge, in units of Newtons?
1
Expert's answer
2018-09-10T12:52:08-0400
According to the superposition principle, resulting force on the charge considered can be derived as a sum of all forces which act on this charge. These forces can be obtained by Coulomb’s law. Calculate the force between +5 C and +7 C charges:
F_1=1/(4πε_0 )·(q_1 q_2)/r^2 =1/(4·3.142·8.854·〖10〗^(-12) )·(5·7)/〖0.2〗^2 =7.864·〖10〗^12 "N."
Calculate the force between +5 C and -1 C charges:
F_2=1/(4πε_0 )·(q_1 q_3)/r^2 =1/(4·3.142·8.854·〖10〗^(-12) )·(5·(-1))/〖1.4〗^2 =-2.293·〖10〗^10 "N."
Therefore,
F=F_1+F_2=7.864·〖10〗^12-2.293·〖10〗^10=7.84·〖10〗^12 N.